The following refers to a purchase of a machine by a company on January 1, 2015.

             Contract Price:                                $80,000                                    Interest Rate:                              6%

             Contract Price Due On:                 December 31, 2017                 Interest Compounded:        semi-annually

The entry to record the purchase of the machine would include a:

A.

debit to machine for $80,000

B.

credit to note payable for $66,960

C.

debit to interest expense for 12,800

D.

debit to discount for $13,040

E.

debit to machine for $122,135

Determine interest expense to be recorded on December 31, 2015.

A.

$2,069

B.

$4,258

C.

$4,258

D.

$2,173

E.

$4,274
Answer #1

Solution 1:

Present value of note = $80,000 * PV factor at 3% for 6th period = $80,000 * 0.837 = $66,960

Discount on notes payable = $80,000 - $66,960 = $13,040

The entry to record the purchase of the machine would include a "debit to discount for $13,040"

Hence option D is correct.

Solution 2:

interest expense to be recorded on December 31, 2015 = $66,960*103% * 3% = $2,069

Hence option A is correct.